in an ideal dual combustioncycle, the pressure , volume, and temperature at the beginning of the adiabatic compression are 93kn/m2 , 0.05m3, and 24 degrees celcius respectively. The volume ratio of the adiabatic compression is 9:1 . The constant volume heat addition pressure ratio is 1.5:1 and the constant pressure heat addition volume ratio is 2:1

Determine , for this cycle;

a. The pressure , volume and temperature at the cycle process change points

b. The thermal efficiency of the cycle

c. The work done per cycle
Answer #1

0 Given data A 2 5 A At initial point ! P, = 93KM 93kPa Temb T, 24+273 m? V = 0.05 m3 - 297K Volume ratio for adiabaticCorspe2 Const pr. heat additions notune ratio (2) V , = 2 2 16) V2 To find in PV, T at all points. for proces 1-2 ire adiabatic ComNO 1 = or 0.05 m3, 5.5555x10 5333 9 rol From eglis Ta T r E تاه 75.2427315 P2 8.4 1.4 297 92 P2 =2015.6840461 kPa for brocers4 LO T3 115.2427315 Il T3 = 1072.864 K volwa los P₂ = 1.5P2 B P2 = 1.5x201 5.6840.46 | Pz - 3023.526069 KRA for process 3-4,5 Ty = 2x1072.1641k Th=2145.728K 2 V₂ V2 V4 = 2X V3 VĄ 2x 0-05 g V2=0.011111m3 for process 4.5. Adiabatic Expansion. 포 equil0.4 2145-724 3023.526 1.4 Ps 1175.641442 = 368.142696 kra 5 by Thermal Efficiency RI ...l s I - 11 r ar(2-1) + (4-1), d=1.5,그 c) Hoek dove per cycle = Head adeled - heat Rejected. Note - Heat addition to take aking place. at const volume é e 2-3 aid& 38.41082 KI At point I pri =MRTI 93 x 0.05 = mx 0.287X 297 m = 0.054552493 kg. betting all dator in eq (3) Hork done Cu(T2-
Answer #2
You’re awesome
source: Funaab
answered by: James